OpenStudy (anonymous):

Solve 3x^2 = –12x – 15

OpenStudy (anonymous):

wouldn't it be 3^2=-12-15??

OpenStudy (anonymous):

Psssst, how far did you get ?

OpenStudy (anonymous):

I haven't started this one yet lol

OpenStudy (anonymous):

x=-2+i,-2-i

OpenStudy (anonymous):

a = 3 b = 12 c = 15? Or a = -3 b = -12 c = -15?

OpenStudy (anonymous):

I solved it and i got a completely off answer.

OpenStudy (anonymous):

good

OpenStudy (anonymous):

first write it as: ax^2+bx+c=0

OpenStudy (anonymous):

I already did. -3x^2 -12x - 15 = 0

OpenStudy (anonymous):

hhhmmmm.. -3x^2 -12x - 15 = 0 is the same as: 3x^2 +12x + 15 = 0 right ?

OpenStudy (anonymous):

Yeah lol

OpenStudy (anonymous):

ok, so a = 3 b = 12 c = 15 now what ?

OpenStudy (anonymous):

what is the discriminant value ?

OpenStudy (anonymous):

\[12 \pm \sqrt{12{^2} -4(3)(15)} \div -6\]

OpenStudy (anonymous):

it's good practice to calculate the discriminant separately first. what do you get for that ?

OpenStudy (anonymous):

I ended up with \[2 \pm 6i\]

OpenStudy (anonymous):

That's not the right answer. It's actually pretty simple if you follow always follow some steps in the same way. But would need your attention for that :)

OpenStudy (anonymous):

Wassup Gorgeous

OpenStudy (anonymous):

still no progress on this problem

OpenStudy (anonymous):

LagrangeSon.. SHH! SmartButLazy.. Hello, stranger. Fiddlearound.. And I no it's not :( I can't figure it out. -.-

OpenStudy (anonymous):

show me how you calculate the discriminant, and show all work :) \[\Delta=(b)^2-4(a)(c)\]

OpenStudy (anonymous):

12^2 - 4(3)(15) 144 - 12(15) 144 - 180 -36

OpenStudy (anonymous):

OK now this: \[x=\frac{-(b)\pm \sqrt{\Delta}}{2(a)}\] Use the draw tool if it's easier. It must be clear how the fraction looks - what's on top and what's on the bottom

OpenStudy (anonymous):

Just plug a, b and the discriminant in the above and show me exactly what you get. no skipping ahead any steps :)

OpenStudy (anonymous):

Is it clear what needs to be done ? If not let me know

OpenStudy (anonymous):

good

OpenStudy (anonymous):

I meant this draw tool though :)

OpenStudy (anonymous):